OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 7 mm and radius of 19 mm? Use 3.14 to approximate pi, and round answers to two decimal places if necessary. ________ mm3

OpenStudy (anonymous):

2646.27 i think

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

btw i already know the formula already

OpenStudy (wolf1728):

cone volume = PI * r^2 * h / 3 cone volume = 3.14 * 19^2 * 7 / 3 cone volume = 2,644.93

OpenStudy (anonymous):

Ya sure?

OpenStudy (wolf1728):

When I make PI = 3.14159265 then I get 2,646.267 The other number was calculated with PI = 3.14

OpenStudy (anonymous):

oh ya its just without the comma on the answer

OpenStudy (wolf1728):

But you are supposed to use PI approximated to 3.14 right?

OpenStudy (anonymous):

A cone-shaped pile of rocks has a diameter of 12 ft and a height of 4 ft. What is the approximate volume of the pile of rocks? Use 3.14 to approximate pi, and round answers to two decimal places if necessary. ________ft^3

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

3.14 * 6^2 * 4 / 3 = 3.14 * 48 = 150.72

OpenStudy (anonymous):

The height of a cone-shaped statue is 9 ft, and the diameter is 4 ft. What is the volume of the statue? Use 3.14 to approximate pi, and round answers to two decimal places if necessary. ________ft3

OpenStudy (anonymous):

Last last question

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Pweez

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

vol = 3.14 * 2^2 * 9 / 3 vol = 3.14 * 12 = 37.68

OpenStudy (anonymous):

YOUR RIGHT!!!!!!

OpenStudy (anonymous):

YA YAY YAY 100% WOHOOOOOOOOO!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

WOOP WOOP!

OpenStudy (anonymous):

-DANCES!- GO WOLF!

OpenStudy (anonymous):

WOOP WOOP!

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Thank you thank you thank you thank you XDDDDDDDDDDDDDDDDDDDDDDDDDDDDD